A random variable X has the probability density function given below

{x ,0<=x< 1

f (x)={2-x, 1<=x<2 Show graphically and find out the robabilities. p(-1<x<55)

{ 0, otherwise


1
Expert's answer
2021-08-22T17:39:39-0400

p(-1<x<55) = F(55) - F(-1)= F(x>2)-F(x<0) = 0